OpenStudy (anonymous):

describe how the graphs of y=x and y=x-15 are related

OpenStudy (anonymous):

a. the graphs have the same shape. the y intercept or y=x is 0, and the x intercept of the second graph is -15. b. the graphs have the same y intercept. the second graph is steeper than y=x. c. the 2 graphs are the same. d. the graphs have the same shape. the y intercept of y=x is 0, and the y intercept of the second graph is -15

jimthompson5910 (jim_thompson5910):

what are the slopes of each?

OpenStudy (anonymous):

what do you mean?

jimthompson5910 (jim_thompson5910):

what is the slope of y = x

jimthompson5910 (jim_thompson5910):

think of it as y = 1x + 0 and compare/match with y = mx+b

OpenStudy (anonymous):

y=x-15 is a translation of y=x. y=x-15 has the same shape as y=x but it is shifted down 15 units

jimthompson5910 (jim_thompson5910):

Jared25 is correct. Both are the same shape so to speak due to the fact that the two slopes are both 1. So they are pointing in the same direction more or less. This means the two lines are parallel.

OpenStudy (anonymous):

so would the answer be c?

jimthompson5910 (jim_thompson5910):

no, jared pointed out that they are NOT the same

jimthompson5910 (jim_thompson5910):

if they were the same, then you'd see the same graph done twice. That means you'd just see 1 line (because one line lies perfectly on top of the other)

OpenStudy (anonymous):

Nope C says that that the are same. while the shape of the graph is the same the are on different spots of the graph

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

im sorry i just learned this today

OpenStudy (anonymous):